Calculate the standard cell potential (E_cell) for:
H2O2(aq) +ClO2(aq) --> ClO2^- + O2(g). (Basic solution). AND
Fe(s) +2Fe^3+(aq) --> 3Fe^2+ (aq)
Answer #1

(a): H2O2(aq) +ClO2(aq) --> ClO2^- + O2(g). (Basic solution).

Oxidation - half cell: H2O2(aq) + 2OH-(aq) ------- > O2(g) + 2H2O + 2e- ; E0​(oxi) = + 1.10 V

Reduction - half cell: 2ClO2(aq) + 2e- ------ >2 ClO2-(aq) ; E0(red) = + 0.95 V

Hence  E0​(cell) =  E0​(oxi) +  E0​(red) = 1.10 V + 0.95 V = 2.05 V (answer)

Note: Use the E value given in your notebook to get exact answer.

(b): Fe(s) +2Fe^3+(aq) --> 3Fe^2+ (aq)

Oxidation - half cell: Fe(s) ------- > Fe2+(aq) + 2e- ; E0​(oxi) = + 0.44 V

Reduction - half cell: 2Fe3+(aq) + 2e- ------ > 2Fe2+(aq) ; E0(red) = + 0.77 V

Hence  E0​(cell) =  E0​(oxi) +  E0​(red) = 0.44 V + 0.77 V = 1.21 V (answer)